OpenStudy (anonymous):

Please help!? .-. Express the following function, F(x) as a composition of two functions f and g. F(x) = x^2 all over (x^2 + 4) I don't even understand what this question is asking? Help me... ?

OpenStudy (anonymous):

you have lots of choices, but the one that springs to mind is \(g(x)=x^2\) and \(f(x)=\frac{x}{x+4}\) so that \[f\circ g(x)=f(x^2)=\frac{x^2}{x^2+4}\]